# A Short History of (Orbital) Decay: Roman’s Prospects for Detecting Dying Planets

### v1

@@ Abstract
software: astropy (Collaboration et al., 2022), numpy (Harris et al., 2020), pandas (McKinney, 2010), matplotlib (Hunter, 2007)
@@ 1 Introduction
It has been known for some time that the excitation of tides by a satellite can result in orbital decay (Counselman, 1973), which is thought to have played a role in the evolution of the Solar System (Goldreich & Soter, 1966). Until the discovery of the first exoplanet around a Sun-like star (Mayor & Queloz, 1995), there was only the Solar System and its one formation outcome to study. Since the first observation of an exoplanetary transit (Charbonneau et al., 2000), surveys have been conducted from the ground (e.g. Alonso et al., 2004; Bakos et al., 2004; Pollacco et al., 2006; Pepper et al., 2007) and space (e.g. Borucki et al., 2010; Howell et al., 2014; Ricker et al., 2015) to detect and characterize transiting planets. These surveys have detected thousands of short- and medium-period planets around hosts with a wide range of stellar properties.
@@
Several studies (e.g. Maciejewski et al., 2018; Patra et al., 2020; Adams et al., 2024) have sought to detect exoplanet orbital decay, but there remain only two robust direct detections: WASP-12 b (Maciejewski et al., 2016; Patra et al., 2017) and Kepler-1658 b (Vissapragada et al., 2022). Indirect evidence suggests that decay sculpts the population of exoplanets we observe. Schlaufman & Winn (2013) argue hot Jupiters are found less frequently around subgiant stars where tidal effects have shorter timescales, though the number of hot Jupiters around red giant stars complicates this picture (Grunblatt et al., 2019). Main sequence hot Jupiter host stars are younger than average, suggesting a hot Jupiter destruction mechanism (Hamer & Schlaufman, 2019; Miyazaki & Masuda, 2023). A recent detection of a transient associated with a planetary engulfment (De et al., 2023) further motivates searches for orbital decay.
@@
The Nancy Grace Roman Space Telescope is NASA’s upcoming astrophysics flagship mission (Spergel et al., 2015). Roman will conduct the Galactic Bulge Time Domain Survey (GBTDS) [^1] https://science.nasa.gov/mission/roman-space-telescope/galactic-bulge-time-domain-survey/ as one of its Core Community Surveys to detect cold and free-floating planets using microlensing (Bennett & Rhie, 2002; Penny et al., 2019; Johnson et al., 2020) and complete a census of planets. The GBTDS will take advantage of Roman’s wide field-of-view and efficiency to observe $\sim 240\times 10^{6}$ stars down to 25 mag in the broad (0.93 – 2.00-micron) F146 filter. High-precision observations towards the Galactic bulge over a long temporal baseline will produce science beyond microlensing, including the detection of an estimated $\sim 100,000$ warm and hot transiting planets (Bennett & Rhie, 2002; Montet et al., 2017; Wilson et al., 2023).
@@
There is substantial historical precedence for the detection of transiting planets from microlensing surveys, including the first planet discovered by transits, OGLE-TR-56 b (Udalski et al., 2002). While radial velocities were used to confirm this planet (Konacki et al., 2003), this will generally not be possible with Roman given the faintness of most of the observed stars.
@@
False positives are ubiquitous in transit surveys (Fressin et al., 2013), and robust vetting will be essential for Roman transiting exoplanet science. Various eclipsing binary (EB) configurations are especially nefarious astrophysical false positives (Cameron, 2012), as they can mimic transit signals. Techniques employed by the Kepler and TESS surveys (e.g., centroid shifts, ellipsoidal variation, secondary eclipse detection, astrodensity profiling, Batalha et al., 2013; Sullivan et al., 2015; Kipping, 2014) could be used for vetting candidates. Chromaticity checks using the transit depth inferred from observations in Roman’s secondary filter(s) could provide an additional means of validating short-period planet candidates.
@@
Roman will expand transiting exoplanet science by 1) discovering planets in unexplored regions of the Galaxy, up to $\sim$20 kpc away, that span the thin disk, thick disk, and bulge and 2) expanding the known transiting planet population by at least an order of magnitude (Wilson et al., 2023). This will enable demographic studies that probe the entire Galactic population of planets. The detection of an unprecedentedly large sample of planets will enable searches for intrinsically rare phenomena, such as orbital decay.
@@
Perhaps the largest uncertainty in estimating a precise orbital decay yield comes from the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, whose estimates can vary by orders of magnitude (e.g. Ogilvie & Lin, 2007) from approximately $10^{5.5}$ to $>\!10^{8}$ (Adams et al., 2024). The value of $Q_{*}^{{}^{\prime}}$ depends on the physical processes by which energy dissipates (Zahn, 2008). Previous work has suggested $Q_{*}^{{}^{\prime}}$ depends on stellar mass, with the decay time only being short on the subgiant branch for the more massive stars (Weinberg et al., 2024), and on the tidal forcing frequency (i.e. the orbital period, see, e.g., Ogilvie & Lin, 2007; Collier Cameron & Jardine, 2018).
@@ 2 Roman and the GBTDS
The GBTDS will be carried out by the Wide Field Instrument (WFI) (Spergel et al., 2015). The technical specifications of the WFI are described in Akeson et al. (2019), and the most up-to-date values are listed on a NASA webpage. [^2] https://roman.gsfc.nasa.gov/science/WFI_technical.html The WFI will leverage Roman’s 2.4-m aperture primary mirror and 18 H4RG-10 4K x 4K detectors with broad near-infrared (NIR) sensitivity to achieve a field of view of 0.281 deg2, $\sim 200$ times larger than Hubble’s WFC3, while maintaining a similar resolution of $0.11^{\prime\prime}$/pixel (Mosby et al., 2020). The WFI instrument includes eight overlapping filters spanning 0.48 – 2.3 micron. The primary F146 survey filter spans 0.93 – 2.00 micron with a PSF FWHM of 0.105 arcsec.
The Survey Definition Committee recently finalized its survey strategy recommendations [^3] https://asd.gsfc.nasa.gov/roman/comm_forum/forum_17/Core_Community_Survey_Reports-rev03-compressed.pdf. These recommendations differ from the notional survey defined in Penny et al. (2019) in numerous ways, though most of these changes are of minimal consequence for orbital decay science. We proceed considering the notional survey, which has the following specifications:
@@
The Survey Definition Committee has laid out three distinct observing scenarios (underguide, nominal, and overguide). The number and placement of fields, season length, and cadence varies for each scenario. Since recent work advocated for changes to the season configuration (e.g. Gould, 2024) and cadence (e.g. Kupfer et al., 2023), we focus on the ramifications of these two survey parameters for detecting orbital decay in Sec. 4.1. We additionally consider the usefulness of extended observations for confirming decay candidates identified in the GBTDS.
Wilson et al. (2023) performed a pixel-level simulation of a single field of the Penny et al. (2019) notional survey to estimate a GBTDS yield of 60,000similar-toabsent60000$\sim$ 60,000∼ 60 , $000200,$ 000 transiting planets. This simulation includes $\sim 59\times 10^{6}$ stars with F146 $<$ 21 as possible transiting planet hosts and led to a catalog of $\sim 9,000$ detected planets in one field. Wilson et al. (2023) followed the approach of Barclay et al. (2018) to simulate the exoplanet population, drawing the number of planets per star from a Poisson distribution consistent with estimated occurrence rates from Hsu et al. (2019) and then assigning each planet its properties and a host star. The planet occurrence rates were not scaled with metallicity, and thus the differences in the metallicities of the host stars in the inner thin disk, thick disk, and bulge, and the impact of metallicity on planet occurrence are neglected. As the occurrence rate of at least some types of planets increases with metallicity (e.g., Gonzalez 1997; Santos et al. 2001; Fischer & Valenti 2005; Buchhave et al. 2014), the yield may be underestimated.
@@ 3.1 Transit Time Modeling
In order to detect orbital decay, a quadratic ephemeris model with a shrinking period must be statistically preferred over the simpler linear ephemeris model. An approximate expression for period decay (Goldreich & Soter, 1966; Adams et al., 2024) is
@@
With the catalog of simulated detected planets from Wilson et al. (2023) and an assumption about the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, we can estimate the rate of orbital decay experienced by each system. We then need to estimate how well Roman will be able to constrain $dP/dt$.
@@ 3.2 Transit Time Covariance Matrix
To estimate the uncertainty with which we can constrain $dP/dt$, we follow the formalism laid out in Gould (2003) and Gould (2024). To make the analytic derivation simpler, we imagine observing one transit epoch at the central time of each season. Since the length of seasons (72 days) is small compared to the survey baseline (4.7 years), this simplification introduces only a small error, as we discuss in Sec. 3.4. We discuss an alternative season configuration in Sec. 4.1. Considering each season as a single observation translates to observing epochs
@@
Because the defined epochs in Eq. 4 are distributed symmetrically about $E=0$, any matrix element $b_{ij}$ where $i+j$ is odd is exactly 00. This leads to an inverse covariance matrix of
@@ 3.3 Analytic Uncertainty
from Carter et al. (2008), where
@@
is the total signal-to-noise ratio of the transit in the limit $r\to 0$, $\theta\equiv\tau/T$ is the ratio of ingress/egress time to the total transit time $T$, $\Gamma$ is the sampling rate or inverse cadence, and $\sigma_{phot}$ is photometric uncertainty per exposure. To first order, $\theta\sim\sqrt{\delta}$ so
@@ 3.4 Validation
where $M$ is the full number of epochs that would actually be observed by Roman. Both of these approximations are good to less than $1\%$ for periods $<5$ days because the observing seasons are short compared to the survey baseline. This suggests we can estimate the uncertainty for each parameter accurately.
@@ 4.1 Survey Optimization and Extensions
Although the GBTDS is essentially defined, we investigate the implications of survey parameters on orbital decay science. Gould (2024) suggested an alternative season configuration ([1, 1, 0, 0, 1, 1, 0, 0, 0, 1, 1], 1 = ON, 2 = OFF) better optimized for black hole science. This configuration translates to observing epochs
@@
using the values from Eq. 22. The Gould configuration leads to an uncertainty in $dP/dt$ that is 70% of the uncertainty in the notional survey. Gould (2024) similarly found this improvement for measurements of astrometric accelerations with Roman.
The effects of other survey parameters are more subtle since they might also appreciably affect the number of transiting planets detected in addition to the uncertainty with which we can measure orbital decay. The cadences under consideration range from 12.1 to 14.8 min, suggesting a modest $\sim 10\%$ effect on measuring uncertainty in $dP/dt$ due to differences in the number of data points. The Survey Definition Committee ultimately recommended $5-7$ primary bulge fields plus a field at the Galactic Center. For large radii, high signal-to-noise ratio detections, this yield is approximately proportional to the number of sources, and thus going from 7 to 5 fields would decrease the yield by $\sim 30\%$. Assuming nothing about the properties of each field, each additional field simply increases the probability of finding a high-significance decay system because there are more systems in total.
Gould et al. (2024) advocated for doubling the sampling rate relative to the baseline Penny et al. (2019) survey design to $\Gamma=8/{\rm hr}$ (or one observation every 7.5 minutes), in order to increase the yield of low-mass free-floating planets detected by microlensing. This would require halving the number of fields, and would result in an increase in the yield of planets with small radii, but a decrease in the yield of planets with larger radii. If there exists a large population of small, short period planets that are just below detectability with the nominal cadences of every 12.4 to 14.8 minutes, then this strategy could increase the yield of tidal decay detections dramatically. However, the Survey Definition Committee ultimately did not consider such high cadences. This may be an interesting strategy to consider in future campaigns during and extended Roman mission.
@@ 4.2 Catalog Detections
We analyzed the catalog of simulated planet detections from Wilson et al. (2023). It provides a planetary mass using the mass-radius relationship of Chen & Kipping (2017), stellar mass, planetary orbital period, stellar radius, planetary radius, transit duration, diluted transit depth, F146 magnitude, and number of transits observed for each simulated detection. It simulated a single field of transiting planet detections in the GBTDS. Fig. 3 shows the 9279 simulated detections in the space of period and radius. We find a few dozen planets within the Roche radii of their host stars and eliminate them from further analysis. We additionally assume eccentricity $e=0$ in all cases.
Figure 3: The distribution of the planets from the Wilson et al. (2023) simulated planet catalog in period and radius. The color bar indicates the apparent magnitude of each planet’s host star.
@@
To estimate the yield of orbital decay detections, we need a model for the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, since there is little consensus on the mechanism or efficiency of tidal dissipation. There are several parametrizations of $Q_{*}^{{}^{\prime}}$ that scale with the planet’s orbital period and occasionally other system properties (e.g Essick & Weinberg, 2016; Penev et al., 2018; Millholland et al., 2025). For our baseline estimate, we assumed a constant $Q_{*}^{{}^{\prime}}=10^{6}$. Fig. 4 shows the resulting distributions of period derivatives $dP/dt$ using Eq. 1 and their uncertainties using Eq. 19.
We then calculated the orbital decay signal-to-noise ratio, SNR $=(dP/dt)/(\sigma_{dP/dt})$ for each planet in the catalog. The system properties of the four highest SNR examples are reported in Table 1. They are massive planets orbiting bright stars with period derivatives generally much larger than those inferred for WASP-12 b and Kepler-1658 b.
@@
Figure 5: Example transit timing for planet #1923 in Table 1 from the GBTDS using the Penny et al. (2019) notional survey. This planet has the most robust detection of orbital decay in the simulated catalog.
Fig. 6 shows the distribution of the orbital decay signal-to-noise ratios $>10^{-3}$ for planets in the catalog using our baseline estimate assuming $Q_{*}^{{}^{\prime}}=10^{6}$. Since Roman will find $\sim 100,000$ transiting planets, we must set a SNR threshold such that it is unlikely random fluctuations produce statistical false positives. SNR $=(dP/dt)/\sigma\simeq\sqrt{\Delta\chi^{2}},$ where $\Delta\chi^{2}$ compares the goodness-of-fit of the linear and quadratic models. $\Delta\chi^{2}$ follows a $\chi^{2}$ distribution with one degree of freedom assuming uncorrelated, Gaussian noise. $P(\Delta\chi^{2}>19.5)\simeq 10^{-5},$ suggesting SNR $\simeq\sqrt{19.5}\simeq 4.4$ is a reasonable threshold for detection. We thus expect that $1-2$ instances of orbital decay would be detectable for this catalog.
Figure 6: The histogram of orbital decay SNRs $>10^{-3}$ using $Q_{*}^{{}^{\prime}}=10^{6}$.
The actual number of bulge fields in the survey will be $5-7$. Since the catalog only simulates the yield for one field, we expect the total yield to be 5$\sim$ 5$\times$ 5 × larger, so we expect $\sim 5-10$ detections across the entire survey.

### v2

@@ Abstract
software: astropy (collaboration_astropy_2022), numpy (harris_array_2020), pandas (mckinney-proc-scipy-2010), matplotlib (hunter_matplotlib_2007)
@@ 1 Introduction
It has been known for some time that the excitation of tides by a satellite can result in orbital decay (counselman_outcomes_1973), which is thought to have played a role in the evolution of the Solar System (goldreich_q_1966). Until the discovery of the first exoplanet around a Sun-like star (mayor_jupiter-mass_1995), there was only the Solar System and its one formation outcome to study. Since the first observation of an exoplanetary transit (charbonneau_detection_2000), surveys have been conducted from the ground (e.g. Alonso_TrES1_2004; bakos_wide-field_2004; pollacco_wasp_2006; pepper_kilodegree_2007) and space (e.g. borucki_kepler_2010; howell_k2_2014; ricker_transiting_2015) to detect and characterize transiting planets. These surveys have detected thousands of short- and medium-period planets around hosts with a wide range of stellar properties.
@@
Several studies (e.g. maciejewski_planet-star_2018; patra_continuing_2020; adams_doomed_2024) have sought to detect exoplanet orbital decay, but there remain only two robust direct detections: WASP-12 b (maciejewski_departure_2016; patra_apparently_2017) and Kepler-1658 b (vissapragada_possible_2022). Indirect evidence suggests that decay sculpts the population of exoplanets we observe. schlaufman_evidence_2013 argue hot Jupiters are found less frequently around subgiant stars where tidal effects have shorter timescales, though the number of hot Jupiters around red giant stars complicates this picture (grunblatt_giant_2019). Main sequence hot Jupiter host stars are younger than average, suggesting a hot Jupiter destruction mechanism (hamer_hot_2019; miyazaki_evidence_2023). A recent detection of a transient associated with a planetary engulfment (de_infrared_2023) further motivates searches for orbital decay.
@@
The Nancy Grace Roman Space Telescope is NASA’s upcoming astrophysics flagship mission (spergel_wide-field_2015). Roman will conduct the Galactic Bulge Time Domain Survey (GBTDS) [^1] https://science.nasa.gov/mission/roman-space-telescope/galactic-bulge-time-domain-survey/ as one of its Core Community Surveys to detect cold and free-floating planets using microlensing (bennett_simulation_2002; penny_predictions_2019; johnson_predictions_2020) and complete a census of planets. The GBTDS will take advantage of Roman’s wide field-of-view and efficiency to observe $\sim 240\times 10^{6}$ stars down to 25 mag in the broad (0.93 – 2.00-micron) F146 filter. High-precision observations towards the Galactic bulge over a long temporal baseline will produce science beyond microlensing, including the detection of an estimated $\sim 100,000$ warm and hot transiting planets (bennett_simulation_2002; montet_measuring_2017; wilson_transiting_2023).
@@
There is substantial historical precedence for the detection of transiting planets from microlensing surveys, including the first planet discovered by transits, OGLE-TR-56 b (udalski_optical_2002). While radial velocities were used to confirm this planet (konacki_extrasolar_2003), this will generally not be possible with Roman given the faintness of most of the observed stars.
@@
False positives are ubiquitous in transit surveys (fressin_false_2013), and robust vetting will be essential for Roman transiting exoplanet science. Various eclipsing binary (EB) configurations are especially nefarious astrophysical false positives (cameron_extrasolar_2012), as they can mimic transit signals. Techniques employed by the Kepler and TESS surveys (e.g., centroid shifts, ellipsoidal variation, secondary eclipse detection, astrodensity profiling, batalha_planetary_2013; sullivan_transiting_2015; kipping_characterizing_2014) could be used for vetting candidates. Chromaticity checks using the transit depth inferred from observations in Roman’s secondary filter(s) could provide an additional means of validating short-period planet candidates.
@@
Roman will expand transiting exoplanet science by 1) discovering planets in unexplored regions of the Galaxy, up to $\sim$20 kpc away, that span the thin disk, thick disk, and bulge and 2) expanding the known transiting planet population by at least an order of magnitude (wilson_transiting_2023). This will enable demographic studies that probe the entire Galactic population of planets. The detection of an unprecedentedly large sample of planets will enable searches for intrinsically rare phenomena, such as orbital decay.
@@
Perhaps the largest uncertainty in estimating a precise orbital decay yield comes from the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, whose estimates can vary by orders of magnitude (e.g. ogilvie_tidal_2007) from approximately $10^{5.5}$ to $10^{8}$ (adams_doomed_2024). The value of $Q_{*}^{{}^{\prime}}$ depends on the physical processes by which energy dissipates (2008EAS....29...67Z). Previous work has suggested $Q_{*}^{{}^{\prime}}$ depends on stellar mass, with the decay time only being short on the subgiant branch for the more massive stars (weinberg_orbital_2024), and on the tidal forcing frequency (i.e. the orbital period, see, e.g., ogilvie_tidal_2007; collier_cameron_hierarchical_2018).
@@ 2 Roman and the GBTDS
The GBTDS will be carried out by the Wide Field Instrument (WFI) (spergel_wide-field_2015). The technical specifications of the WFI are described in akeson_wide_2019, and the most up-to-date values are listed on a NASA webpage. [^2] https://roman.gsfc.nasa.gov/science/WFI_technical.html The WFI will leverage Roman’s 2.4-m aperture primary mirror and 18 H4RG-10 4K x 4K detectors with broad near-infrared (NIR) sensitivity to achieve a field of view of 0.281 deg2, $\sim 200$ times larger than Hubble’s WFC3, while maintaining a similar resolution of $0.11^{\prime\prime}$/pixel (mosby_properties_2020). The WFI instrument includes eight overlapping filters spanning 0.48 – 2.3 micron. The primary F146 survey filter spans 0.93 – 2.00 micron with a PSF FWHM of 0.105 arcsec.
The Roman Observations Time Allocation Committee provided their recommendations for the strategies for the Core Community Surveys [^3] https://roman.gsfc.nasa.gov/science/ccs/ROTAC-Report-20250424-v1.pdf. These recommendations differ from the notional survey defined in penny_predictions_2019 in numerous ways, though most of these changes are of minimal consequence for orbital decay science. We proceed considering the notional survey, which has the following specifications:
@@
The Survey Definition Committee has laid out three distinct observing scenarios (underguide, nominal, and overguide). The number and placement of fields, season length, and cadence varies for each scenario. Since recent work advocated for changes to the season configuration (e.g. gould_one_2024) and cadence (e.g. kupfer_continuous_2023), we focus on the ramifications of these two survey parameters for detecting orbital decay in Sec. 4.1. We additionally consider the usefulness of extended observations for confirming decay candidates identified in the GBTDS.
wilson_transiting_2023 performed a pixel-level simulation of a single field of the penny_predictions_2019 notional survey to estimate a GBTDS yield of $\sim 60,000$$200,000$ transiting planets. This simulation includes $\sim 59\times 10^{6}$ stars with F146 << 21 as possible transiting planet hosts and led to a catalog of $\sim 9,000$ detected planets in one field. wilson_transiting_2023 followed the approach of barclay_revised_2018 to simulate the exoplanet population, drawing the number of planets per star from a Poisson distribution consistent with estimated occurrence rates from hsu_occurrence_2019 and then assigning each planet its properties and a host star. Only main-sequence and subgiant stars with $\log{g}\geq 3.5$ were considered as viable planet hosts.
The catalog we use did not scale planet occurrence rates with metallicity, whose relationship is well-studied (e.g., Gonzalez:1997; Santos:2001; Fischer:2005; Buchhave:2014), albeit not well characterized for all planet populations. When accounting for the higher metallicities of the inner Galaxy, wilson_transiting_2023 found overall yield increases by a factor of $\sim 3$, the yield of $4-8\,R_{\oplus}$ planets increases by $\sim 5$, and the yield of $8\,R_{\oplus}$ planets increases by $\sim 2$. The detections of orbital decay we find are of giant planets, and thus if the scaling of planet frequency with metallicity adopted by wilson_transiting_2023 holds for the stellar populations probed by Roman, we can expect a yield of decaying planets that is a few times larger than our baseline predictions.
@@ 3.1 Transit Time Modeling
In order to detect orbital decay, a quadratic ephemeris model with a shrinking period must be statistically preferred over the simpler linear ephemeris model. An approximate expression for period decay (goldreich_q_1966; adams_doomed_2024) is
@@
With the catalog of simulated detected planets from wilson_transiting_2023 and an assumption about the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, we can estimate the rate of orbital decay experienced by each system. We then need to estimate how well Roman will be able to constrain $dP/dt$.
@@ 3.2 Transit Time Covariance Matrix
To estimate the uncertainty with which we can constrain $dP/dt$, we follow the formalism laid out in gould_chi2_2003 and gould_one_2024. To make the analytic derivation simpler, we imagine observing one transit epoch at the central time of each season. Since the length of seasons (72 days) is small compared to the survey baseline (4.7 years), this simplification introduces only a small error, as we discuss in Sec. 3.4. We discuss an alternative season configuration in Sec. 4.1. Considering each season as a single observation translates to observing epochs
@@
Because the defined epochs in Eq. 4 are distributed symmetrically about $E=0$, any matrix element $b_{ij}$ where $i+j$ is odd is exactly 0. This leads to an inverse covariance matrix of
@@ 3.3 Analytic Uncertainty
from carter_analytic_2008, where
@@
is the total signal-to-noise ratio of the transit in the limit $r\to 0$, $\theta\equiv\tau/T$ is the ratio of ingress/egress time to the total transit time $T$, $\Gamma$ is the sampling rate or inverse cadence, and $\sigma_{phot}$ is photometric uncertainty per exposure. Roman’s $\sigma_{phot}$ is shown in Fig. 4 of wilson_transiting_2023, and for stars with an F146 mag $\sim 18$, $\sigma_{phot}\approx 0.2\%$ To first order, $\theta\sim\sqrt{\delta}$ so
@@ 3.4 Validation
where $M$ is the full number of epochs that would actually be observed by Roman. Both of these approximations are good to less than $1\%$ for periods <5<5 days because the observing seasons are short compared to the survey baseline. This suggests we can estimate the uncertainty for each parameter accurately.
@@ 4.1 Survey Optimization and Extensions
Although the GBTDS survey strategy is essentially defined, we investigate the implications of survey parameters on orbital decay science. gould_one_2024 suggested an alternative season configuration ([1, 1, 0, 0, 1, 1, 0, 0, 0, 1, 1], 1 = ON, 2 = OFF) better optimized for black hole science. This configuration translates to observing epochs
@@
using the values from Eq. 22. The Gould configuration leads to an uncertainty in $dP/dt$ that is 70% of the uncertainty in the notional survey. gould_one_2024 similarly found this improvement for measurements of astrometric accelerations with Roman.
The effects of other survey parameters are more subtle since they might also appreciably affect the number of transiting planets detected in addition to the uncertainty with which we can measure orbital decay. The cadences under consideration range from 12.1 to 14.8 min, suggesting a modest $\sim 10\%$ effect on measuring uncertainty in $dP/dt$ due to differences in the cadence at fixed exposure time. However, the decreased cadence results in an increase in the exposure time at fixed overhead time, leading to a further improvement in the uncertainty in $dP/dT$. The Survey Definition Committee ultimately recommended 575-7 primary bulge fields plus a field at the Galactic Center. For large radii, high signal-to-noise ratio detections, this yield is approximately proportional to the number of sources, and thus going from 7 to 5 fields would decrease the yield by $\sim 30\%$. Assuming nothing about the properties of each field, each additional field simply increases the probability of finding a high-significance decay system because there are more systems in total.
2024arXiv240614531G advocated for doubling the sampling rate relative to the baseline Penny et al. (2019) survey design to $\Gamma=8/{\rm hr}$ (or one observation every 7.5 minutes), in order to increase the yield of low-mass free-floating planets detected by microlensing. This would require halving the number of fields, and would result in an increase in the yield of planets with small radii, but a decrease in the yield of planets with larger radii. If there exists a large population of small, short-period planets that are just below detectability with the nominal cadences of every 12.4 to 14.8 minutes, then this strategy could increase the yield of tidal decay detections dramatically. However, the Survey Definition Committee ultimately did not consider such high cadences. This may be an interesting strategy to consider in future campaigns during an extended Roman mission.
@@ 4.2 Catalog Detections
We analyzed the catalog of simulated planet detections from wilson_transiting_2023. It provides a planetary mass using the mass-radius relationship of chen_probabilistic_2017, stellar mass, planetary orbital period, stellar radius, planetary radius, transit duration, diluted transit depth, F146 magnitude, and number of transits observed for each simulated detection. It simulated a single field of transiting planet detections in the GBTDS. Fig. 3 shows the 9279 simulated detections in the space of period and radius. We find a few dozen planets within the Roche radii of their host stars and eliminate them from further analysis. We additionally assume eccentricity $e=0$ in all cases.
Figure 3: The distribution of the planets from the wilson_transiting_2023 simulated planet catalog in period and radius. The color bar indicates the apparent magnitude of each planet’s host star.
@@
To estimate the yield of orbital decay detections, we need a model for the stellar tidal dissipation factor $Q_{*}^{{}^{\prime}}$, since there is little consensus on the mechanism or efficiency of tidal dissipation. There are several parametrizations of $Q_{*}^{{}^{\prime}}$ that scale with the planet’s orbital period and occasionally other system properties (e.g essick_orbital_2016; penev_empirical_2018; millholland_empirical_2025). For our baseline estimate, we assumed a constant $Q_{*}^{{}^{\prime}}=10^{6}$. Fig. 4 shows the resulting distributions of period derivatives $dP/dt$ using Eq. 1 and their uncertainties using Eq. 19.
@@
We then calculated the orbital decay signal-to-noise ratio, SNR $(dP/dt)/(\sigma_{dP/dt})$ for each planet in the catalog. The system properties of the four highest SNR examples are reported in Table 1. They are massive planets orbiting bright stars with period derivatives generally much larger than those inferred for WASP-12 b and Kepler-1658 b.
Figure 5: Example transit timing for planet #1923 in Table 1 from the GBTDS using the penny_predictions_2019 notional survey. This planet has the most robust detection of orbital decay in the simulated catalog.
@@
Figure 6: The histogram of orbital decay SNRs $10^{-3}$ using three distinct values of $Q_{*}^{{}^{\prime}}$.
Fig. 6 shows the distribution of the orbital decay signal-to-noise ratios $10^{-3}$ for planets in the catalog using our baseline estimate assuming $Q_{*}^{{}^{\prime}}=10^{6}$ and alternative scenarios $Q_{*}^{{}^{\prime}}=10^{5}$ and $Q_{*}^{{}^{\prime}}=10^{7}$. Since Roman will find $\sim 100,000$ transiting planets, we must set a SNR threshold such that it is unlikely random fluctuations produce statistical false positives. SNR $(dP/dt)/\sigma\simeq\sqrt{\Delta\chi^{2}},$ where $\Delta\chi^{2}$ compares the goodness-of-fit of the linear and quadratic models. $\Delta\chi^{2}$ follows a $\chi^{2}$ distribution with one degree of freedom assuming uncorrelated, Gaussian noise. $P(\Delta\chi^{2}>19.5)\simeq 10^{-5},$ suggesting SNR $\simeq\sqrt{19.5}\simeq 4.4$ is a reasonable threshold for detection. We thus expect that 121-2 instances of orbital decay would be detectable for this catalog assuming $Q_{*}^{{}^{\prime}}=10^{6}$.
Two detections out of the $\sim 3600$ hot Jupiters (P<10P<10\,days and $R_{p}>5\,R_{\oplus}$) in the catalog is a fraction of $\sim 6\times 10^{-4}$, which is modestly smaller than the currently inferred fraction of decay systems due to the faintness of Roman stars. As shown in Fig. 6, decay is highly sensitive to the value of $Q_{*}^{{}^{\prime}}$, as the catalog yield goes from 0 to $2$ to $\sim 30$ when $Q_{*}^{{}^{\prime}}$ goes from $10^{7}$ to $10^{6}$ to $10^{5}$.
The actual number of bulge fields in the survey will be $5$. Since the catalog only simulates the yield for one field, we expect the total yield to be $\sim 5\times$ larger, our baseline estimate is that there will be $\sim 5-10$ instances of orbital decay detectable across the entire survey.